How much of an alloy that is 30% copper should be mixed with 200 ounces of an alloy that is 90% copper in order to get an alloy that is 70% copper?
Answer by scott8148(6628)

x(30%)+(200)(90%)=(200+x)(70%) ... .3x+180=140+.7x ... 40=.4x ... 100=x